# Are LLMs Prescient? A Continuous Evaluation using Daily News as the Oracle

### v1

@@ Abstract
Many existing evaluation benchmarks for Large Language Models (LLMs) quickly become outdated due to the emergence of new models and training data. These benchmarks also fall short in assessing how LLM performance changes over time, as they consist of static questions without a temporal dimension. To address these limitations, we propose using future event prediction as a continuous evaluation method to assess LLMs’ temporal generalization and forecasting abilities. Our benchmark, Daily Oracle, automatically generates question-answer (QA) pairs from daily news, challenging LLMs to predict “future” event outcomes. Our findings reveal that as pre-training data becomes outdated, LLM performance degrades over time. While Retrieval Augmented Generation (RAG) has the potential to enhance prediction accuracy, the performance degradation pattern persists, highlighting the need for continuous model updates.
@@ 1 Introduction
Traditional Large Language Model (LLM) benchmarks are often static, and do not reflect real-world information that evolves over time. This presents two significant challenges. First, as LLMs are updated, there is a risk that the static benchmarks become outdated and more vulnerable to data leakage, where their content might end up in the training data of newer models. This undermines the reliability of performance assessments on these benchmarks (Sainz et al.,, 2023; Xu et al.,, 2024; McIntosh et al.,, 2024; Li and Flanigan,, 2024). Second, the static benchmarks often lack the temporal information to track the model’s performance variations over time (McIntosh et al.,, 2024). This creates a need for evaluation methods that stay relevant over time and incorporate temporal dynamics.
Daily news provides a natural setting for continuous evaluation of LLMs. Since the world is constantly changing, a benchmark designed around forecasting the next day’s news will never be out of date by construction. In addition to enabling continuous evaluation, forecasting is itself a longstanding challenge with significant implications across various domains, including healthcare, finance, and policymaking (Tetlock and Gardner,, 2016; Dempsey et al.,, 2017; Gillingham et al.,, 2018; Lopez-Lira and Tang,, 2023). While human experts have traditionally made such forecasts, machine learning models, particularly LLMs, have emerged as promising alternatives due to their capability to learn from vast and diverse corpora (Halawi et al.,, 2024; Ye et al.,, 2024; Yan et al.,, 2023). Several recent forecasting question-answer (QA) datasets have been developed (Jin et al.,, 2021; Zou et al.,, 2022; Zhang et al.,, 2024), however, they are limited in either size, scope, or do not continuously keep pace with the rapidly changing world. More critically, the extent to which LLMs’ predictive abilities change over time remains underexplored.
@@
We continuously evaluate various LLMs, both with and without access to a limited archive of news articles. Our experiments reveal that LLMs experience an average performance decline of 20.14% on True/False (TF) questions and 23.26% on Multiple Choice (MC) questions between January 2020 and September 2024, with degradation becoming more pronounced before and after the models’ knowledge cutoff dates. Although models utilizing Retrieval Augmented Generation (RAG) (Lewis et al.,, 2020) can demonstrate improved prediction performance, the downward trend persists, suggesting an ongoing challenge in maintaining temporal generalization. Overall, our benchmark highlights the challenges posed by outdated pertaining data in LLMs, and underscores the necessity for continuous model updating to keep up with the constantly evolving stream of real-world information.
@@
  • Empirical Findings on Performance Degradation: Our experiments demonstrate a consistent decline in LLM prediction accuracy over time, even when supplemented with recent “open-book” information. The persistent degradation highlights the challenges posed by outdated pretraining data in LLMs and underscores the necessity for continuous model updating.
@@ 2 Related Work
Lazaridou et al., (2021) define temporal generalization as the ability of Language Models to generalize well to future data from beyond their training period. They demonstrate that Transformer-XL’s performance deteriorates over time, evidenced by increasing perplexity when evaluated on post-training data. However, perplexity-based metrics have two main limitations: they cannot be applied to closed-source models lacking accessible logits, and increased perplexity does not necessarily indicate degraded performance on downstream tasks (Röttger and Pierrehumbert,, 2021; Agarwal and Nenkova,, 2022). Zhu et al., (2024) investigate temporal generation using the Bits Per Character (BPC) metric. Similar to perplexity, BPC fails to capture higher-level performance on downstream tasks. In contrast, our work focuses on evaluating how well models acquire and utilize real-world event knowledge in downstream forecasting tasks, providing a more nuanced assessment of temporal generalization.
@@
While static QA datasets evaluate models on fixed knowledge snapshots, dynamic QA datasets incorporate a temporal dimension, allowing assessment of how models adapt to evolving information. Several dynamic QA datasets are proposed. Chen et al., (2021) construct TimeQA by using time-sensitive facts in WikiData with aligned Wikipedia passages to synthesize QA pairs. Zhang and Choi, (2021) introduce SituatedQA by manually annotating temporally and geographically dependent questions. StreamingQA (Liska et al.,, 2022) and RealtimeQA (Kasai et al.,, 2024) are both dynamic benchmarks with QA pairs answerable from news articles. StreamingQA, however, does not provide continuous evaluation with always-relevant data. RealTimeQA does not address forecasting and is more like a plugin for a search engine, in the sense that it tests whether a model has updated its knowledge as facts change, rather than testing whether it can predict what will change given its knowledge of the past. FreshQA (Vu et al.,, 2024) contains a fixed set of human-written open-ended questions whose answers by nature can change based on new developments in the world, but is smaller and does not address forecasting. It is also updated weekly rather than daily. While all these datasets have some form of time-sensitivity like the Daily Oracle, they either do not provide continuous evaluation or do not evaluate forecasting capabilities, or neither.
@@
Forecasting questions aim to assess a model’s ability to predict the outcomes of future events based on its existing knowledge. Several datasets in the event forecasting field have been introduced. ForecastQA (Jin et al.,, 2021) used crowdworkers to collect 10,392 QA pairs from news articles. Zou et al., (2022) argue that the QA pairs from ForecastQA are often nonsensical or ambiguous since they are written by humans without forecasting expertise. They further introduce AutoCast, a forecasting dataset from popular human forecasting tournaments containing 6,707 QA pairs. In contrast, our Daily Oracle dataset is generated automatically from daily news articles, which means that it is never out of date, can easily grow its size automatically without additional inputs from human forecasters, and provides more comprehensive event coverage than human forecasting tournaments.
Similar to our generation method, TLB-forecast (Zhang et al.,, 2024) has an automatic forecasting QA generation framework using news articles. However, their dataset is constrained both temporally and topically, and only contains cooperation and conflict events in Middle-Eastern countries from 2015 to 2022. This restricts the dataset from evaluating more general event-prediction abilities. Furthermore, considering most of the powerful LLMs have been developed after 2020, the portions of the dataset covering earlier years may contain answers already seen during training. This prior exposure compromises the dataset’s effectiveness as a forecasting benchmark. In contrast, our dataset spans a broader timeframe and covers more topics, offering a more comprehensive and reliable forecasting benchmark.
Note that none of the aforementioned datasets provide insights into how prediction ability changes over time. Zhu et al., (2024) introduce FreshBench, a forecasting dataset scraped from the GoodJudgmentOpen platform, and also study temporal generalization. However, they report accuracy in a relatively short time window (from January 2023 to April 2024) with only 2,532 questions. While we observe a gradual performance decline in our dataset, they report significant fluctuations in model accuracy shortly after release. We argue that our dataset presents a more challenging task, as evidenced by the continuous degradation in model performance over time, making it a robust benchmark for assessing LLMs’ temporal generalization.
@@
Dataset Continuous? Interval Forecast? Size Latest Update
TimeQA None 20,000 2021
SituatedQA None 4,757 2021
StreamingQA None 36,800 2021
RealTimeQA None 1,470 2023
FreshQA Weekly 600 2024
ForecastQA None 10,382 2019
AutoCast None 6,707 2022
TLB-forecast None 6,604 2022
FreshBench Unknown 2,532 2024
Daily Oracle (Ours) Daily 29,988 2024
Table 1: We compare Daily Oracle with existing benchmarks in the literature. For continuously updated datasets (e.g. Daily Oracle, FreshBench, and FreshQA), “Interval” refers to the dataset update interval, and “Size” and “Latest Update” refer to the fixed data currently available. Our Daily Oracle benchmark is the only forecasting benchmark which is continuously updated every day using questions generated from daily news.
@@ 3 The Daily Oracle Dataset
In this section, we present Daily Oracle, a continuously updated QA benchmark of forecasting questions that are automatically generated from daily news. For our current analysis of LLM performance, we utilize a subset of the data consisting of 16,082 TF questions and 13,906 MC questions, covering a diverse range of forecasting topics, which are generated using daily news articles from January 2020 up until September 2024. However, our QA generation framework is continuous and updates daily. In section 3.1, we describe our LLM-based dataset construction pipeline, detailing the data sources and the four-step construction process. Section 3.2 provides an analysis and general overview of the dataset.
@@ 3.1 Dataset Construction
Figure 1: Data Construction Process of Daily Oracle.
@@
Following Zou et al., (2022), we collect a large corpus of news articles from the daily-updated Common Crawl News Dataset (Nagel,, 2016) with the news-please package (Hamborg et al.,, 2017). We filter for mainstream sources—CBS News, CNBC, CNN, Forbes, and NPR. While our data collection and evaluation are performed daily, for this study we utilize a static news corpus with 1,216,925 English articles spanning January 2019 to September 2024. This corpus is also used for the constrained open-book evaluation setting in section 4.1.
@@
QA pairs are generated from articles published between January 2020 and September 2024.111For news articles in 2019, we use them as the corpus for the constrained open-book setting. Due to budget constraints, for each day, we select six articles for QA generation: three are chosen randomly, and three are selected from hot topics. Details for how hot topics are chosen can be found in Appendix A.1. For each selected article, we then use LLM to generate two TF QA pairs and two MC QA pairs with the few-shot prompting technique. [^2] See Appendix C for all the prompts we use.
Inspired by Zhang et al., (2024), our QA construction follows four steps, as illustrated in Figure 1:
@@
    QA Filtering. In the final step, we prompt the LLM to check seven principles: correctness of answers, non-answerability before the publication date, absence of information leakage, objectivity, inclusion of a clear temporal element, public interest, and non-obviousness of the answer. Each principle is scored with 0, 1, or 2 points, and we selected the questions that received at least 13 points in total.
These principles are detailed in Appendix A.2. We use GPT-3.5 (OpenAI,, 2024) for steps (1) and (4), while GPT-4 (OpenAI,, 2023) is utilized for steps (2) and (3) to ensure high data quality, as GPT-4 is currently the best-performing LLM available.
@@ 3.2 Dataset Analysis
At the time of writing this paper, the subset dataset we use from Daily Oracle consists of 16,082 TF and 13,906 MC QA pairs, covering the period from January 1st, 2020, to September 30th, 2024, with an average of 17.3 questions per day. Figure 2(a) shows that our dataset covers various MC question types, mainly starting with “What will” (26.9%), “Who will” (21.0%), and “Which … will” (18.4%). Figure 2(b) provides a breakdown of the categories, highlighting our dataset’s broad coverage. The categorization of each question is determined using GPT-3.5, based on the prompt from Halawi et al., (2024). Examples of QA pairs are shown in Table 2.
(a) MC question types for Daily Oracle.
(b) Category Breakdown for Daily Oracle.
Figure 2: Comparison of question types and categories.
@@
TF Politics & Governance Will the prosecution’s key witness in the New York hush money trial in April 2024 be someone other than Michael Cohen? –No.
TF Politics & Governance Will the House Energy and Commerce Committee vote unanimously to advance a bill that could potentially ban TikTok if ByteDance does not sell the app by March 2024? –Yes.
MC What Science & Tech What will be the starting price range for the Google Pixel 8a as of May 2024? A.\$599–\$649 B. \$199–\$249 C. \$750–\$800, D. \$499–\$559. –D.
MC Who Sports Who will go on the injured list before the New York Mets’ game on May 29, 2024? A. Pete Alonso B. Edwin Diaz C. Jeff McNeil D. Francisco Lindor –B.
MC Which Arts & Recreation By May 2024, on which streaming service will “The First Omen” become available for subscribers? A. Disney+, B. Hulu, C. Amazon Prime Video, D. Netflix –B.
MC How many Science & Tech How many U.S. states will the path of totality cross during the total solar eclipse on April 8, as reported by February 2024? A. 15 B. 10 C. 20 D. 6 –A.
MC Where Healthcare & Biology Where will the second known U.S. case of bird flu in a human be reported by March 2024? A. California, B. Texas, C. New York, D. Florida –B.
MC How much Economics & Business How much will Apple, Inc. (AAPL) be up year-to-date by the end of June 2024? A. Up 149.5% B. Just over 19% C. 9.7%. D. 27%. –C.
Table 2: Daily Oracle Example Questions and Answers.
@@
Each question in Daily Oracle implicitly requires the model to retrieve relevant knowledge. How do these requirements change day by day over the course of our benchmark? Anderson and Schooler, (1991) study a similar relationship in human information environments, specifically in New York Times headlines, children’s verbal interactions, and emails. In Figure 3, we take inspiration from their work and analyze whether a word’s frequency of occurrence in the past 100 days predicts its occurrence on the next day. In other words, if over the past 100 days we have frequently required the model to retrieve specific knowledge, e.g. if there are many questions about the unemployment rate, is it likely it will have to retrieve this knowledge in the future?
We analyze this relationship for words in the titles of the articles we use to generate questions as well as in the text of the TF and MC questions themselves. Past frequency is computed by checking, for each day in the 100 day window, if a word has occurred in any article title (so, the maximum frequency is 100). We find that there is a linear relationship between the frequency of usage in the past 100 days and the probability of occurrence on the 101st day in all cases, replicating Anderson & Schooler’s findings for New York Times headlines. Interestingly, there is a drop in probability for both TF and MC questions for words with a frequency of 40, though it is unclear why. There is also some clustering at lower frequencies, particularly in the TF and MC question plots. Many words appear less than 20 times during the 100 day window. The temporal structure exhibited in the daily news stream may be of a future point of interest from a modeling perspective.
Figure 3: Following Anderson and Schooler, (1991), we plot the probability of a word occurring in an (left) article title, (middle) True/False question, or (right) Multiple Choice question given how frequently it had appeared in one over the past 100 days, computed over our entire dataset. We fit a linear regression in each case and show a linear relationship in each case ($R^{2}=0.978$, $0.986$, and $0.985$ for left, middle, and right respectively).

### v2

@@ Abstract
Many existing evaluation benchmarks for Large Language Models (LLMs) quickly become outdated due to the emergence of new models and training data. These benchmarks also fall short in assessing how LLM performance changes over time, as they consist of a static set of questions without a temporal dimension. To address these limitations, we propose using future event prediction as a continuous evaluation method to assess LLMs’ temporal generalization and forecasting abilities. Our benchmark, Daily Oracle, automatically generates question-answer (QA) pairs from daily news, challenging LLMs to predict “future” event outcomes. Our findings reveal that as pre-training data becomes outdated, LLM performance degrades over time. While Retrieval Augmented Generation (RAG) has the potential to enhance prediction accuracy, the performance degradation pattern persists, highlighting the need for continuous model updates. Code and data are available at https://agenticlearning.ai/daily-oracle.
Continuous Evaluation, News Forecasting
@@ 1 Introduction
Traditional Large Language Model (LLM) benchmarks are often static, and do not reflect real-world information that evolves over time. This presents two significant challenges. First, as LLMs are updated, there is a risk that static benchmarks become outdated and more vulnerable to data leakage, where their content might end up in the training data of newer models. This undermines the reliability of performance assessments on these benchmarks (Sainz et al., 2023; Xu et al., 2024; McIntosh et al., 2025; Li & Flanigan, 2024). Second, static benchmarks often lack temporal information, making it difficult to track models’ performance variations over time (McIntosh et al., 2025). This creates a need for evaluation methods that always remain relevant and incorporate temporal dynamics.
Daily news provides a natural setting for continuous evaluation of LLMs. Since the world is constantly changing, a benchmark designed around forecasting the next day’s news will never be out of date by construction. In addition to enabling continuous evaluation, forecasting is itself a longstanding challenge with significant implications across various domains, including healthcare, finance, and policymaking (Tetlock & Gardner, 2016; Dempsey et al., 2017; Gillingham et al., 2018; Lopez-Lira & Tang, 2023). While human experts have traditionally made such forecasts, machine learning models, particularly LLMs, have emerged as promising alternatives due to their ability to learn from vast and diverse corpora (Halawi et al., 2024; Ye et al., 2024; Yan et al., 2024). Several recent forecasting question-answer (QA) datasets have been developed (Jin et al., 2021; Zou et al., 2022; Zhang et al., 2024), however, they are limited in either size, scope, or they do not continuously keep pace with the rapidly changing world. More critically, the extent to which LLMs’ predictive abilities change over time remains understudied.
@@
We continuously evaluate various LLMs, both with and without access to a limited archive of news articles. Our experiments reveal that LLMs experience significant performance degradation between January 2020 and December 2024, with degradation becoming more pronounced before and after the models’ knowledge cutoff dates. On average, performance drops by 21.55% on TF questions and 11.33% on MC questions. While model performance can be improved with more recent news articles using Retrieval Augmented Generation (RAG) (Lewis et al., 2020), the downward trend persists, suggesting the challenge in maintaining its prediction ability over time.
@@
  • Empirical Findings on Performance Degradation: Since our benchmark provides new questions each day, we can study how model performance changes along the temporal axis. Our work effectively reveals a clear performance degradation pattern in LLMs’ forecasting accuracy over time. Additionally, we study how this pattern changes as the LLMs are given access to updated knowledge up to different times. Surprisingly, we find that, even when the model has access to recent information in an “open-book” setting, it still experiences performance degradation. Moreover, the sheer degree of decline, along with its smoothness over time, was unexpected. On the one hand, this highlights the problems with outdated LLM pre-training data and on the other hand underscores the need for continuous model updating.
@@ 2 Related Work
Dataset Continuous? Interval Forecast? Size Latest Update
TimeQA (Chen et al., 2021) None 20,000 2021
SituatedQA (Zhang & Choi, 2021) None 4,757 2021
StreamingQA (Liska et al., 2022) None 36,800 2021
RealTimeQA (Kasai et al., 2024) None 1,470 2023
FreshQA (Vu et al., 2024) Weekly 600 2024
ForecastQA (Jin et al., 2021) None 10,382 2019
AutoCast (Zou et al., 2022) None 6,707 2022
ForecastBench (Karger et al., 2025) Biweekly 1,000 2024
TLB-forecast (Zhang et al., 2024) None 6,604 2022
FreshBench (Zhu et al., 2025) Unknown 2,769 2024
Daily Oracle (Ours) Daily 31,510 2024*
    Our experiments use the subset generated until December 2024. Daily Oracle itself remains active, continuing to generate new questions daily from 2025 onward.
Table 1: We compare Daily Oracle with existing benchmarks in the literature. For continuously updated datasets (e.g. Daily Oracle, FreshQA, FreshBench, and ForecastBench), “Interval” refers to the dataset update interval, and “Size” and “Latest Update” refer to the fixed data currently available. Our Daily Oracle benchmark is the only forecasting benchmark which is continuously updated every day using questions generated from daily news.
@@
Lazaridou et al. (2021) define temporal generalization as the ability of Language Models to generalize well to future data from beyond their training period. They demonstrate that Transformer-XL’s performance deteriorates over time, evidenced by increasing perplexity when evaluated on post-training data. However, perplexity-based metrics have two main limitations: they cannot be applied to closed-source models lacking accessible logits, and increased perplexity does not necessarily indicate degraded performance on downstream tasks (Röttger & Pierrehumbert, 2021; Agarwal & Nenkova, 2022). Zhu et al. (2025) investigate temporal generalization using the Bits Per Character (BPC) metric. Similar to perplexity, BPC fails to capture higher-level performance on downstream tasks. In contrast, our work focuses on the downstream forecasting task, evaluating how well models understand world knowledge and make predictions. This approach offers a more reliable evaluation of temporal generalization with direct relevance to real-world applications and public interest.
@@
While static QA datasets evaluate models on fixed knowledge snapshots, dynamic QA datasets incorporate a temporal dimension, allowing assessment of how models adapt to evolving information. Several dynamic QA datasets are proposed. Chen et al. (2021) construct TimeQA by using time-sensitive facts in WikiData with aligned Wikipedia passages to synthesize QA pairs. Zhang & Choi (2021) introduce SituatedQA by manually annotating temporally and geographically dependent questions. StreamingQA (Liska et al., 2022) and RealtimeQA (Kasai et al., 2024) are both dynamic benchmarks with QA pairs answerable from news articles. StreamingQA, however, does not provide continuous evaluation with always-relevant data. RealTimeQA does not address forecasting and is more like a plugin for a search engine, in the sense that it tests whether a model has updated its knowledge as facts change, rather than testing whether it can predict what will change given its knowledge of the past. FreshQA (Vu et al., 2024) contains a fixed set of human-written open-ended questions whose answers by nature can change based on new developments in the world, but is smaller and does not address forecasting. It is also updated weekly rather than daily. While all these datasets have some form of time-sensitivity like the Daily Oracle, they either do not provide continuous evaluation or do not evaluate forecasting capabilities, or neither.
@@
Forecasting questions aim to assess a model’s ability to predict the outcomes of future events based on its existing knowledge. Several datasets in the event forecasting field have been introduced. ForecastQA (Jin et al., 2021) used crowdworkers to collect 10,392 QA pairs from news articles. Zou et al. (2022) argue that the QA pairs from ForecastQA are often nonsensical or ambiguous since they are written by humans without forecasting expertise. They further introduce AutoCast, a forecasting dataset from popular human forecasting tournaments containing 6,707 QA pairs. While ForecastQA and AutoCast remain static, ForecastBench (Karger et al., 2025) regularly updates a set of 1,000 forecasting questions either sourced from forecasting markets or generated via fixed templates based on real-world event datasets. However, it still depends on users actively submitting new forecasting questions or maintaining the underlying datasets. In contrast, our Daily Oracle dataset is generated automatically from daily news articles, which means that it is never out of date, can easily grow its size automatically without additional inputs from human forecasters, and provides more comprehensive event coverage than human forecasting tournaments.
Similar to our generation method, TLB-forecast (Zhang et al., 2024) has an automatic forecasting QA generation framework using news articles. However, their dataset is constrained both temporally and topically, only containing cooperation and conflict events in Middle-Eastern countries from 2015 to 2022. This restricts the dataset from evaluating more general event-prediction abilities. Furthermore, considering most of the powerful LLMs have been developed after 2020, the portions of the dataset covering earlier years may contain answers already seen during training. This prior exposure compromises the dataset’s effectiveness as a forecasting benchmark. In contrast, our dataset spans a broader timeframe and covers more topics, offering a more comprehensive forecasting benchmark.
Note that none of the aforementioned datasets provide insights into how prediction ability changes over time. Zhu et al. (2025) introduce FreshBench, a forecasting dataset scraped from the Good Judgment Open platform, and also study temporal generalization. However, they report accuracy in a relatively short time window (from January 2023 to August 2024) with only 2,769 questions. While we observe a gradual performance decline in our dataset, they report significant fluctuations in model accuracy shortly after release. A closer look reveals key limitations of forecasting market–based questions for studying temporal generalization: they suffer from limited early coverage, inconsistent distribution over time, and reduced dataset size after filtering due to a high proportion of low-quality questions, making it difficult to reliably analyze temporal performance trends. In contrast, our automatically generated dataset has broad event coverage, consistent growth, and more uniform question quality over time. [^1] See Appendix C for details on comparing LLM-generated and forecasting market datasets.
@@ 3 The Daily Oracle Dataset
In this section, we present Daily Oracle, a continuously updated QA benchmark of forecasting questions that are automatically generated from daily news. For our current analysis of LLM performance, we utilize a subset of the data consisting of 16,783 TF questions and 14,727 MC questions, covering a diverse range of forecasting topics, which are generated using daily news articles from January 2020 up until December 2024. However, our QA generation framework is continuous and updates daily. In Section 3.1, we describe our LLM-based dataset construction pipeline, detailing the data sources and the four-step construction process. Section 3.2 provides an analysis and general overview of the dataset. Lastly, in Section 3.3, we conduct a human evaluation, similar to our QA filtering process, to verify the quality of the generated QA pairs.
@@ 3.1 Dataset Construction
Following Zou et al. (2022), we collect a large corpus of news articles from the daily-updated Common Crawl News Dataset (Nagel, 2016) with the news-please package (Hamborg et al., 2017). To further enrich our news dataset, we supplement it with daily scraped news using the Newspaper3k package. [^2] https://newspaper.readthedocs.io/en/latest/ We filter for mainstream sources—CBS News, CNBC, CNN, Forbes, and NPR. While our data collection and evaluation are performed daily, for this study we utilize a static news corpus with 1,246,973 English articles spanning January 2019 to December 2024. This corpus is also used for the constrained open-book evaluation setting in Section 4.1.
@@
QA pairs are generated from articles published between January 2020 and December 2024.333For news articles in 2019, we use them as the corpus for the constrained open-book setting. For each day, we select six articles for QA generation: three are chosen randomly, and three are selected from hot topics. [^4] See Appendix A.2 for hot-topic selection details. For each selected article, we then use LLM to generate two TF QA pairs and two MC QA pairs with the few-shot prompting technique. [^5] See Appendix D for all the prompts we use.
We adopt the methodology of Zhang et al. (2024), as their prompt design largely suits our setting. To further filter the QA pairs, we establish seven key criteria to ensure they qualify as valid forecasting questions, incorporating these into a QA Filtering step. The QA construction follows four steps, as illustrated in Figure 7:
@@
    QA Filtering. We prompt the LLM to check seven principles: correctness of answers, non-answerability before the publication date, absence of information leakage, objectivity, inclusion of a clear temporal element, public interest, and non-obviousness of the answer. Each principle is scored with 0, 1, or 2 points, and we selected the questions that received at least 13 points in total. These principles are detailed in Appendix A.3.
We use GPT-3.5 (OpenAI, 2024a) or GPT-4o-mini (OpenAI, 2024b) for steps (1) and (4), while GPT-4 (OpenAI, 2023) or GPT-4o (OpenAI, 2024b) is utilized for steps (2) and (3) to ensure high data quality. [^6] We use GPT-3.5 and GPT-4 until September 2024. After October 2024, we switch to GPT-4o-mini and GPT-4o, as they are both more cost-effective and more powerful.
Figure 1: Pie charts showing (a) MC question type distribution and (b) question category distribution in Daily Oracle.
Figure 2: Following Anderson & Schooler (1991), we plot the probability of a word occurring in an (left) article title, (middle) True/False question, or (right) Multiple Choice question given how frequently it had appeared in one over the past 100 days, computed over our entire dataset. We fit a linear regression and show a linear relationship in each case ($R^{2}=0.843$, $0.982$, and $0.990$ for left, middle, and right respectively).
@@ 3.2 Dataset Analysis
At the time of writing this paper, the subset dataset we use from Daily Oracle consists of 16,783 TF and 14,727 MC QA pairs, covering the period from January 1st, 2020, to December 31st, 2024, with an average of 17.2 questions per day. Figure 1(a) shows that our dataset covers various MC question types, mainly starting with “What will” (27.3%), “Who will” (21.2%), and “Which … will” (18.4%). Figure 1(b) provides a breakdown of the categories, highlighting our dataset’s broad coverage. The categorization of each question is determined using GPT-3.5, based on the prompt from Halawi et al. (2024). Examples of QA pairs are shown in Table 2.
Table 2: Daily Oracle Example Questions and Answers.
@@
TF Politics & Governance – Will the prosecution’s key witness in the New York hush money trial in April 2024 be someone other than Michael Cohen? –No.
TF Politics & Governance – Will the House Energy and Commerce Committee vote unanimously to advance a bill that could potentially ban TikTok if ByteDance does not sell the app by March 2024? –Yes.
MC What Science & Tech – What will be the starting price range for the Google Pixel 8a as of May 2024? A.\$599–\$649 B. \$199–\$249 C. \$750–\$800, D. \$499–\$559. –D.
MC Who Sports – Who will go on the injured list before the New York Mets’ game on May 29, 2024? A. Pete Alonso B. Edwin Diaz C. Jeff McNeil D. Francisco Lindor –B.
MC Which Arts & Recreation – By May 2024, on which streaming service will “The First Omen” become available for subscribers? A. Disney+, B. Hulu, C. Amazon Prime Video, D. Netflix –B.
MC How many Science & Tech – How many U.S. states will the path of totality cross during the total solar eclipse on April 8, as reported by February 2024? A. 15 B. 10 C. 20 D. 6 –A.
MC Where Healthcare & Biology – Where will the second known U.S. case of bird flu in a human be reported by March 2024? A. California, B. Texas, C. New York, D. Florida –B.
MC How much Economics & Business – How much will Apple, Inc. (AAPL) be up year-to-date by the end of June 2024? A. Up 149.5% B. Just over 19% C. 9.7%. D. 27%. –C.
@@
Each question in Daily Oracle implicitly requires the model to retrieve relevant knowledge. How do these requirements change day by day over the course of our benchmark? Anderson & Schooler (1991) explored similar patterns in human information environments. Inspired by their work, Figure 2 examines whether a word’s frequency over the past 100 days predicts its occurrence the next day—e.g., if many questions concern the unemployment rate, will this trend continue?
We analyze this relationship for words in the titles of the articles we use to generate questions as well as in the text of the TF and MC questions themselves. Past frequency is computed by checking, for each day in the 100 day window, if a word has occurred in any article title (so, the maximum frequency is 100). We find that there is a linear relationship between the frequency of usage in the past 100 days and the probability of occurrence on the 101st day in all cases, replicating Anderson & Schooler’s findings for New York Times headlines. This indicates that past information usage strongly predicts future retrieval needs, suggesting a temporal structure in the knowledge demands of our benchmark.
3.3 Human Evaluation
To assess the quality of our LLM-based filtering method, we randomly sample 30 TF and 30 MC QA pairs and ask 4 human annotators to evaluate them using the seven principles outlined in the QA Filtering step in Section 3.1. We evaluate the consistency among annotators using Fleiss’ Kappa (Fleiss, 1971), which yields an average inter-rater agreement score of 0.26, indicating fair agreement. We then compute the human consensus score as the average of human scores and compare it to LLM-assigned scores, finding an average accuracy of 89.52% across the 7 principles. For final QA pair acceptance (i.e., threshold above 13 points), the LLM and human consensus scores demonstrated an accuracy of 85.00%, further supporting the reliability of our LLM-based filtering approach. A detailed breakdown of human evaluation metrics is provided in Appendix A.4.